what is the point of intersection of the following straight lines:
y = -9x - 10
-y = -8x + 2

Answer by tarungehlot(9)

add both equation
we gets
0= -17x - 8
x= -8/17
y= 72/17 -10= -98/17
hence point of intersection
(x,y)= ( -8/17,-98/17)